What is the resultant pH of a solution when 20ml of 0.1N NaOH is mixed with 20ml of 0.05M Ca(OH) 2 at 25C?

Grade:10

2 Answers

Saood
5 years ago

when 20ml of 0.1N NaOH is mixed with 20ml of 0.05M Ca(OH)2 at 25C

Equivalent weight of NaOH = molecular weight of NaOH because its acidity is one

NORMALITY of NaOH = MOLARITY of NaOH =.1

Milli mole of NaOH=molarity x volume = 20x.1 = 2

NaOH——->Na+ + OH-

2 milli mole of NaOH produce 2 milli mole of OH-

Milli mole of Ca(OH)2=molarity x volume = 20x.05 = 1

Ca(OH)2———→Ca+2 + 2OH-

1 milli mole of Ca(OH)2 produce 2 milli mole of OH-

Total milli mole of OH- = 1+2 = 3

Total volume of solution(ml)=20+20=40ml

OH- CONCENTRATION =Milli mole of OH-/Total volume of solution(ml)=3/40M

pOH=-Log[OH-] = -log3/40 = log40-log3 =1.6020 -.4771 = 1.1249

pH=14–1.1249=12.8751

SAMBHAV MISHRA
5 years ago
@saood u idiot u have copied the answer from quora.. if u dont have the knowledge then dont answer.. or else i will complain against you idiot